Circular No. 8328

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 H2 (LARSEN) J. A. Larsen, Lunar and Planetary Laboratory, reports his discovery

of a comet (discovery observation below) on Spacewatch images taken with the 0.9-m telescope at Kitt Peak; the object showed a coma diameter of 7′′ and a 15′′ tail in p.a. 240^{o}. Following posting on the NEO Confirmation Page, other CCD observers also noted the object’s cometary appearance, including J. G. Ries (McDonald Observatory, 0.76-m reflector; Apr. 21.24– 21.43 UT; object extended with a hint of tail to the southwest), G. Hug (Scranton, KS, 0.3-m reflector; Apr. 21.4; diffuse), and J. Young (Table Mountain, 0.6-m reflector; Apr. 21.42–21.45; small central condensation in a round 4′′ coma, with a hint of elongation ∼ 8′′ long in p.a. 290^{o}).

2004 UT  α_{2000}  δ_{2000}  Mag.
Apr. 19.41414  16^{h}15^{m}51.^{s}78  −3^{o}22′09′′.8  20.4

The available astrometry, the following preliminary (rather uncertain) elliptical orbital elements, and an ephemeris appear on MPEC 2004-H42.

T = 2004 Aug. 11.946 TT  ω = 132.^{o}957
e = 0.42552  Ω = 128.984 }2000.0
q = 2.58236 AU  i = 12.085
a = 4.49509 AU n^{o} = 0.103418 P = 9.5 years

V2574 OPHIUCHI E. Mason, A. Ederoclite, L. Germany, and F. Salas, European Southern

Observatory, Santiago; and M. Della Valle, Istituto Nazionale di Astrofisica, Arcetri/Firenze, report that échelle spectra (390-900 nm; resolution 48000) of V2574 Oph (cf. IAUC 8323), obtained on Apr. 17.32 and 18.37 UT at La Silla with the 2.2-m telescope (+ Fiber-fed Extended Range Optical Spectrograph), are dominated by Hα, Na D, and Ca II(2) emission lines, which are flanked by double P-Cyg profiles. From the minima of the P-Cyg absorptions, the measured expansion velocities are ∼ 400 and 1050 km/s for Hα and 400 and 1000 km/s for Na D. Emission lines of Fe II (42, 49, and 74) and O I (multiplet 1) were also detected. The overall appearance of the spectra indicates a substantial evolution with respect to the description provided on IAUC 8324, suggesting that this is a slow nova caught at maximum or slightly before maximum light.

Magnitude estimates: Apr. 17.436 UT, V = 10.53 (J. D. West, Mulvane, KS; CCD); 20.882, 10.2 (A. Pearce, Nedlands, W. Australia; visual).

2004 April 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green